# A One Dimensional (1D) Computational Fluid Dynamics Study of Fontan-Associated Liver Disease (FALD)

### v1

@@ 4 Discussion
In this study, we construct a detailed model of the liver vasculature to explore the hemodynamics of patients with HLHS and DORV. We also impose FALD conditions on the HLHS patient to predict pressure and flow under FALD progression. Results show that the HLHS patient has higher pressure than the DORV patient in the portal and hepatic networks. These values reach hypertensive ranges when FALD is imposed. WSS is also higher for HLHS patients, also increasing with FALD severity. The increased WSS is an indicator of vascular remodeling and contributes to the development of fibrosis [61]. The perfusion and pressure distribution plots show the increased pressure changes with increasing levels of FALD, giving insight into regions of tissue that will be affected by vascular remodeling and fibrosis development due to this increase in pressure.
@@ 4.1 Flow and Pressure
Pressure and flow predictions are generated in both hepatic and portal networks. This is a unique contribution of this study in that it considers both inlet networks. The patient parameters are chosen based on a study by Taylor-LaPole et al. [36], and the patient input waveforms are generated using data from the above research.
Model predictions produce a higher average portal venous and hepatic artery pressure for the HLHS patient, as shown in Tables 4 and 5. Although there is no significant variation in the average pressure, there is a significant increase in the predicted systolic and pulse pressure when FALD is imposed, and these values enter hypertensive conditions [62]. The increase in pulse pressure is an indicator of increased arterial stiffening [63]. Portal hypertension and arterial stiffening is a key characteristic of FALD, often associated with the development of fibrosis [11]. Although many patients do not present outward symptoms of FALD until an irreversible stage (i.e., organ failure, liver transplant required), the model is designed to capture the degree of portal hypertension within Fontan patients. Portal hypertension indicates vasculature remodeling, fibrotic development, and venous congestion. It is considered a multiorgan disorder associated with abnormal hemodynamics in the systemic and portal vasculatures [64].
There is no significant change in pressure and flow under FALD conditions in the hepatic artery. This may be caused by the hepatic arterial buffer response (HABR), an intrinsic regulation of hepatic arterial blood flow [65]. Our result agrees with the study by [66], suggesting that HABR is preserved even in patients with advanced cirrhosis with significant portal hypertension. It is also important to note that our model predictions for the hepatic network are consistent with those from the full body network of the same DORV and HLHS patients in the study by Taylor-LaPole et al. [36].
Figure 6: Perfusion plots distributing the average flow in terminal vessels to the liver. Results are shown at two angles (Posterior and Anterior). For each view, the left column depicts the network and average flow in the terminal vessel, and the right column distributes the flow at the end of each terminal vessel to the nearby tissue. Since the inflow to the DORV and HLHS networks remain unchanged with disease, the flow distribution does not vary significantly between cases. In general, the flow decreases slightly as disease severity is increased. Note the light red shown in the anterior view refers to values from 4.8-5.0 mL/s whereas the deep red shown in the posterior view refers to values from 0-0.2 mL/s.
@@ 4.2 Wall Shear Stress
The HLHS patient has higher WSS values than the DORV patient, as shown in Table 6. The higher WSS may indicate increased fibrosis and vascular remodeling [61, 67]. As discussed in Taylor-LaPole et al., increased WSS in the HLHS patient is not uncommon, as patients with reconstructed aortas tend to have increased WSS values downstream of surgical reconstruction due to vascular remodeling [36, 68]. Our results for the hepatic network can be confirmed by the study by Taylor-LaPole et al. [69], where the WSS for single-ventricle patients continues to increase with decreasing radius. The results in the portal network closely resemble the results obtained by Bartolo et al. [70] in the pulmonary vasculature. Although the portal vein is systemic, its pressure values and physiology closely resemble the arteries within the pulmonary vasculature. A study by Voges et al. [68] found that patients with reconstructive aortic surgery will have downstream vessels try to dilate in the years following surgery to compensate for the decrease in cardiac output. This dilation increases the WSS, which is consistent with the findings in our analysis.
Our predictions show that WSS decreases slightly as the severity of FALD increases. The decrease in WSS is more prominent in smaller downstream vessels. Decreased WSS from a patient’s norm is indicative of increased hypertension and continued stiffening of the vessel wall. An experimental study by Traub et al. [71] found that low WSS is correlated with the upregulation of vasoconstrictive genes. This upregulation leads to increased vascular modeling, fibrosis, and hypertension [72].
@@ 4.3 Perfusion
The perfusion plots reveal that the DORV patient has a more uniform flow distribution throughout the liver tissue than the HLHS patient. For the healthy patient with HLHS, most blood goes to the upper and lower regions of the liver. As FALD progresses, flow in these regions is reduced while flow increases in the upper-center of the liver (yellow/green region). These differences in blood delivery are confirmed with increasing KL divergence values as FALD conditions increase. This is indicative of vascular remodeling in small downstream vessels [73]. It is important to note that blood flow changes occur in tissue regions fed by the portal network. This change in flow, most likely the result of vascular remodeling, is a direct response to portal hypertension as the body tries to relieve downstream resistance in regions of the liver that develop fibrosis [74]. Although this change in flow can reveal indicators of FALD progression, it is not routine practice to monitor blood flow within the liver until patients develop severe outward symptoms. CT and ultrasound are most commonly used to monitor the development of fibrosis [24]. This demonstrates the importance of the mathematical model from which blood distribution in the liver can be detected to provide clinicians with information if more investigation or intervention is required.
The most common indicator of advanced FALD is portal hypertension [24]. The plots showing the pressure distribution demonstrate the pressure differences in specific tissue regions between patient types. The DORV patient has the most consistent pressure distribution throughout the tissue, with very slight differences in the pressures in the tissues fed by the portal network. The pressure in these portal regions increases in the healthy HLHS patient. The left posterior side of the liver has an increase in pressure, and the upper center portion of the anterior side. These regions continue to increase in pressure, reaching hypertensive values as FALD severity increases. As mentioned above, portal hypertension leads to fibrosis. Fibrosis increases downstream resistance, creating a cycle of continual increasing pressures and reduced blood flow [24].
@@ 5 Limitations
Since our model is one-dimensional, we cannot predict the hemodynamics on different sides of the portal vein and hepatic artery walls. This is relevant because sclerosis could occur, especially when the WSS is low in one area of the vessel [75]. We plan to incorporate a 3D-CFD model in the large portal vein and hepatic artery vessels with our hemodynamic data and liver model in future studies. A description of energy loss can be included in the 1D model, as was done in several previous studies [47, 76]. As discussed by Taylor-LaPole et al. [36], we can incorporate an energy loss model that could be calibrated by comparing 1D and 3D simulations. For example, using a 3D fluid-structure interaction (FSI) model applied to the reconstructed aorta [77], or by identifying potential areas of secondary flow directly from Four-Dimensional Magnetic Resonance Imaging (4D-MRI) data.
Another limitation is the closed-loop nature of this model. We have limited our study to the liver vasculature, neglecting the upstream network that begins at the heart. As a result, the model does not account for the potential hemodynamic disturbances that would propagate from the reconstructed aorta. We used a prescribed inflow waveform predicted from the study by Taylor-LaPole et al. [36]. We did not combine the portal vein and hepatic networks, which may cause some information to be lost. In future studies, we plan to combine the hepatic artery and portal vein networks with a whole-body network containing the patient’s aorta. Additionally, the perfusion analysis was done by mapping predictions at the terminals of vessels to segmented liver tissue. We plan to incorporate a porous media model at the vessel terminals to This will to allow us to account for mixing and drainage of the blood from the two systems. This will give a more complete picture of what is happening within the tissue. We also plan to do more studies were we reduce the cardiac output in accordance with the level of FALD. As FALD progresses and portal hypertension increases, patients will experience further reduced cardiac output. In our study we kept cardiac output consistent across all HLHS and FALD patients. Insight from our clinical collaborators along with increased access to data will make this future study possible.
During calibration, we manually adjusted the stiffness parameters (see Table 2) to fit an expected pressure curve and the portal venous pressure of a healthy individual. However, patients with HLHS are likely to have portal vein hypertension. Thus, the actual pressure of the patient with HLHS may be higher than our predicted pressures. To enhance our model calibration’s precision, we can also incorporate multiple pressure readings from various vessels or utilize multiple blood pressure cuff measurements.
@@ 6 Conclusion
This study uses a patient-specific 1D-CFD model to predict hemodynamics in the portal vein and hepatic artery vasculatures of an HLHS patient with a reconstructed aorta and a DORV patient with a native aorta. Our model was calibrated to the cardiac output and liver inflow data from a previous study on this data set by [36] with parameters in Table 2 agreeing with normal portal and hepatic venous pressures. We projected cardiac output from DORV and HLHS patients onto the liver model extracted from an adult liver CT and used the 1D-CFD model to predict flow and pressure in the liver networks. We simulated FALD conditions in the HLHS patient to predict pressure and flow across various stages of FALD progression. Our results show that the HLHS patient has increased pressure and reduced flow in the portal vein compared to the DORV patient. These results were more prominent as FALD conditions increased. The approximate WSS was lower in both the portal and hepatic vessels in the patient with HLHS, suggesting the presence of hypertension. Perfusion analysis shows that progression of FALD results in increased pressures in large regions of liver tissue, possibly leading to fibrosis development, and reduced flow to upper and lower regions of the tissue. In general, our study offers information on FALD pathology by showing specific changes in the hemodynamics of the Fontan circulation that could cause portal hypertension and complications of FALD. More clinical 4D-MRI studies are recommended to gain deeper insights into the dynamic hemodynamics of FALD and Fontan-associated circulatory systems.

### v2

@@ 4 Discussion
In this study, we construct a detailed model of the liver vasculature to explore the hemodynamics of patients with HLHS and DORV. We also impose FALD conditions on the HLHS patient to predict pressure and flow under FALD progression. The impact of cardiac output and inflow waveform is also investigated by imposing patient inflows on their counterparts. Results show that the HLHS patient has higher pressure than the DORV patient in the portal and hepatic networks. These values reach hypertensive ranges when FALD is imposed. The HLHS network continues to have increased pressures regardless of waveform and parameters imposed. WSS is also higher for HLHS patients in the hepatic network, also increasing with FALD severity. The increased WSS is an indicator of vascular remodeling and contributes to the development of fibrosis [Eshtehard2017]. The perfusion and pressure distribution plots show the increased pressure changes with increasing levels of FALD, giving insight into regions of tissue that will be affected by vascular remodeling and fibrosis development due to this increase in pressure.
@@ 4.1 Flow and Pressure
Pressure and flow predictions are generated in both hepatic and portal networks. This is a unique contribution of this study in that it considers both inlet networks. The patient parameters are chosen based on a study by Taylor-LaPole et al. [alyssa2022], and the patient input waveforms are generated using data from the above research.
Model predictions produce a higher average hepatic artery pressure for the HLHS patient, as shown in Table 5. Although there is only slight variation in the average pressure, there is a significant increase in the predicted systolic and pulse pressure when FALD is imposed, and these values enter potentially hypertensive conditions [banerjee2012portalhypertension]. This increased pressure is observed in the hepatic network, regardless of inflow or parameters imposed. Progression of FALD also causes and increase in pulse pressure in the portal network. The increase in pulse pressure is an indicator of increased vascular stiffening [MayoClinicPulsePressure]. Portal hypertension and vascular stiffening is a key characteristic of FALD, often associated with the development of fibrosis [emamaullee2020fontan]. Although many patients do not present outward symptoms of FALD until an irreversible stage (i.e., organ failure, liver transplant required), the model is designed to capture the degree of hypertension within Fontan patients. Hypertension indicates vasculature remodeling, fibrotic development, and venous congestion. It is considered a multiorgan disorder associated with abnormal hemodynamics in the systemic and portal vasculatures [Macmathuna2012]. Both DORV and HLHS portal pressures increases when their counterpart’s inflow was imposed on their network (Table 4). This could be due to the larger impact of inflow on vessels of smaller radius which is expected by Poiseuille’s law [Poiseuillelaw]. It is also important to note that our model predictions for the hepatic network are consistent with those from the full body network of the same DORV and HLHS patients in the study by Taylor-LaPole et al. [alyssa2022].
Figure 8: Perfusion plots distributing the average flow in terminal vessels to the liver. Results are shown at two angles (Posterior and Anterior). For each view, the left column depicts the network and average flow in the terminal vessel, and the right column distributes the flow at the end of each terminal vessel to the nearby tissue. Since the inflow to the HLHS network remains unchanged with disease, the flow distribution does not vary significantly between HLHS cases. In general, the flow decreases slightly as disease severity is increased. Note the light red shown in the anterior view refers to values from 4.8-5.0 mL/s whereas the deep red shown in the posterior view refers to values from 0-0.2 mL/s.
@@ 4.2 Wall Shear Stress
The HLHS patient has higher WSS values than the DORV patient in the hepatic artery and portal vein, as shown in Table 6. In the hepatic network, WSS increases under FALD conditions. The higher WSS may indicate increased fibrosis and vascular remodeling [Eshtehard2017, Piva2012]. As discussed in [Espina], WSS values in the hepatic artery can range from 1 to 9.5 Pa depending on the location of the arteries they are obtained in. However, these values normally fall within 3 to 8 in the hepatic artery itself and from 1-3 Pa in the downstream hepatic vessels. All HLHS patients have increased hepatic artery WSS values and those with FALD conditions imposed have increased WSS values in the downstream vasculature. The results in the portal network closely resemble the results obtained by Bartolo et al. [bartolo2022numerical] in the pulmonary vasculature. Although the portal vein is systemic, its pressure values and physiology closely resemble the arteries within the pulmonary vasculature. In the portal network, the DORV tends to have higher WSS values in the downstream vasculature. Low WSS values in the portal vein tend to indicate portal hypertension [wei2017wall] and portal hypertension is indicative of liver fibrosis [SukKim]. An experimental study by Traub et al. [traub] found that low WSS is correlated with the upregulation of vasoconstrictive genes. This upregulation leads to increased vascular modeling, fibrosis, and hypertension [iwakiri2014pathophysiology]. While all the WSS values are low compared to the typical value Wei et al. [wei2017wall] defines (1.0 Pa), in downstream vasculature the HLHS patients do have further decreased values compared to the DORV patient.
@@ 4.3 Perfusion
The perfusion plots reveal that the DORV patient has a more uniform flow distribution throughout the liver tissue than the HLHS patient. For the healthy patient with HLHS, most blood goes to the upper and lower regions of the liver. As FALD progresses, flow in these regions is reduced while flow increases in the upper-center of the liver (yellow/green region). These differences in blood delivery are confirmed with increasing KL divergence values as FALD conditions increase. This is indicative of vascular remodeling in small downstream vessels [Kolega_remodel]. It is important to note that blood flow changes occur in tissue regions fed by the portal network. This change in flow, most likely the result of vascular remodeling, is a direct response to portal hypertension as the body tries to relieve downstream resistance in regions of the liver that develop fibrosis [Lemoinne2015]. Although this change in flow can reveal indicators of FALD progression, it is not routine practice to monitor blood flow within the liver until patients develop severe outward symptoms. CT and ultrasound are most commonly used to monitor the development of fibrosis [gordonwalker2019fontanliver]. This demonstrates the importance of the mathematical model from which blood distribution in the liver can be detected to provide clinicians with information if more investigation or intervention is required.
The most common indicator of advanced FALD is portal hypertension [gordonwalker2019fontanliver]. The plots showing the pressure distribution demonstrate the pressure differences in specific tissue regions between patient types. The DORV patient has the most consistent pressure distribution throughout the tissue, with very slight differences in the pressures in the tissues fed by the portal network. The pressure in these portal regions increases in the healthy HLHS patient. The left posterior side of the liver has an increase in pressure, and the upper center portion of the anterior side. These regions continue to increase in pressure, reaching hypertensive values as FALD severity increases. As mentioned above, portal hypertension leads to fibrosis. Fibrosis increases downstream resistance, creating a cycle of continual increasing pressures and reduced blood flow [gordonwalker2019fontanliver].
@@ 5 Limitations
Since our model is one-dimensional, we cannot predict the hemodynamics on different sides of the portal vein and hepatic artery walls. This is relevant because sclerosis could occur, especially when the WSS is low in one area of the vessel [zhou2023wallshearstress]. In future studies, a description of energy loss can be included in the 1D model, as in several previous studies [Colebank21, mynard2015unified]. As Taylor-LaPole et al. [alyssa2022] discussed, we can incorporate an energy loss model that could be calibrated by comparing 1D and 3D simulations, for example, by identifying potential areas of secondary flow directly from Four-Dimensional Magnetic Resonance Imaging (4D-MRI) data.
This study focused on the effects transmitted by the aorta, although clearly the system is connected, and therefore, we should also have added a model of the venous system. Moreover, respiration also affects flow [respiration], another feature not studied here. Although these factors likely influence the findings, they also add complexity to the system, which, without more data, can be challenging to validate.
Another limitation is the network itself. We have limited our study to the liver vasculature, neglecting the upstream network that begins at the heart. As a result, the model does not account for the potential hemodynamic disturbances that would propagate from the reconstructed aorta. We used a prescribed inflow waveform predicted from the study by Taylor-LaPole et al. [alyssa2022]. We did not combine the portal vein and hepatic networks, which can cause some information to be lost. A two-sided model was not incorporated, neglecting hemodynamic effects of the Fontan conduit where some backward flow tends to occur. In future studies, we plan to combine the hepatic artery and portal vein networks with a whole-body, two-sided network containing the patient’s aorta and Fontan conduit. In addition, perfusion analysis was performed by mapping the predictions at the vessels’ terminals to segmented liver tissue. We plan to incorporate a porous media model at the vessel terminals to allow us to account for mixing and drainage of the blood from the two systems. This will give a more complete picture of what is happening within the tissue. We also plan to do more studies to reduce cardiac output according to the level of FALD. As FALD progresses and portal hypertension increases, patients will experience a further reduced cardiac output. In our study, we kept cardiac output consistent across all HLHS and FALD patients. Insight from our clinical collaborators, along with greater access to data, will make this future study possible.
During calibration, we manually adjusted the stiffness parameters (see Table 2) to fit the expected pressure curve and the portal venous pressure of a healthy individual. However, patients with HLHS are likely to have portal vein hypertension. Thus, the actual pressure of the patient with HLHS may be higher than our predicted pressures. To enhance our model calibration’s precision, we can also incorporate multiple pressure readings from various vessels or utilize multiple blood pressure cuff measurements.
@@ 6 Conclusion
This study uses a patient-specific 1D-CFD model to predict hemodynamics in the portal vein and hepatic artery vasculatures of an HLHS patient with a reconstructed aorta and a DORV patient with a native aorta. Our model was calibrated to the cardiac output and liver inflow data from a previous study on this data set by [alyssa2022] with parameters in Table 2 agreeing with normal portal and hepatic venous pressures. We projected cardiac output from DORV and HLHS patients onto the liver model extracted from an adult liver CT and used the 1D-CFD model to predict flow and pressure in the liver networks. We simulated FALD conditions in the HLHS patient to predict pressure and flow across various stages of FALD progression. Our results show that the HLHS patient has increased pressure in the hepatic network compared to the DORV patient, regardless of inflow or parameters imposed. These results were more prominent as FALD conditions increased. The approximate WSS was higher in the hepatic vessels in simulated FALD patients. However, WSS values were lower in the portal network for all HLHS patients. Perfusion analysis shows that progression of FALD results in increased pressures in large regions of liver tissue, possibly leading to fibrosis development, and reduced flow to upper and lower regions of the tissue. In general, our study offers information on FALD pathology by showing specific changes in the hemodynamics of the Fontan circulation that could cause portal hypertension and complications of FALD. More clinical 4D-MRI studies are recommended to gain deeper insights into the dynamic hemodynamics of FALD and Fontan-associated circulatory systems.